A student took a test composed of 20 questions, with correct answers worth 2 points. If the student received 38 points for his grade, how many questions were correct?
Answer by richwmiller(17219)

20*2=40 perfect score
40-38=2 missed two points
2/2=1 missed answer 2 points missed by 2 points per question
20-1=19 correct answers
check
19*2=38
ok